How can I rewrite the expression using parentheses to get the given value of
28-3x3+4 with the value being 23?

Answer by bucky(2189)

I notice that your mathematical expression uses "x" as meaning multiplication.
.
You can set this up as an equation. When you do it looks like this:
.
28 - 3 x 3 + 4 = 23
.
You can make this an equality (both sides are equal) by inserting one set of parentheses as follows:
.
28 - (3 x 3) + 4 = 23
.
Inside of the parentheses the 3 x 3 equals 9. Substituting 9 for the contents inside the parentheses changes the left side to:
.
28 - 9 + 4 = 23
.
If you algebraically total the three numbers on the left side by subtracting 9 from 28 and then adding 4 to that answer, the equation becomes:
.
23 = 23
.
So the answer to your problem is to insert parentheses as shown below and the result will equal 23:
.
28 - (3 x 3) + 4
.
Hope this helps you understand the problem better.